OpenStudy (anonymous):

Jon and Kate are having a pizza party. A large pizza costs $15.25 and small costs $8.25. Jon and Kate purchase 23 pizza's for a total cost of $294.75. How many of each type of pizza did they buy?

OpenStudy (anonymous):

PLEASE HELP :)

OpenStudy (anonymous):

Solve the following for T and s:\[\left\{\frac{1525}{100}L +\frac{825}{100}s=\frac{29475}{100}, \text{ T}+s=23\right\} \]\[\{L=15,s=8\} \]